OpenStudy (anonymous):

make h the subject of the formulae here T=2pi√ H-h/g , please show the steps

OpenStudy (jack1):

hey @ubah just to clarigy: is this ur eqn: \(\Large T=2\pi \sqrt {H-\frac hg}\) ?

OpenStudy (anonymous):

\[T = 2\pi \sqrt{\frac{ H-h }{ g}}\] this is what i meant

OpenStudy (jack1):

ah, cool T=2pi√ H-h/g \(\Large T=2\pi \sqrt { \frac {H-h}g}\) \(\Large \frac T{2\pi}= \sqrt { \frac {H-h}g}\) \(\Large (\frac T{2\pi})^2= (\sqrt { \frac {H-h}g})^2\) \(\Large \frac {T^2}{4\pi^2}= \frac {H-h}g\) \(\Large \frac {T^2 \times g}{4\pi^2}= H-h\)

OpenStudy (jack1):

... i think...?

OpenStudy (anonymous):

am waiting for the final step

OpenStudy (jack1):

well... u could do some of it if u want ;P all that's left is subtract H from both sides... then flip the signs on both sides so you're left with h = ...something... can u do this?

OpenStudy (anonymous):

no @jack1

OpenStudy (alekos):

they all wanna be spoon fed the answers................go figure

OpenStudy (jack1):

sorry @ubah ... it's against OS policy to provide answers, only help HOW to get answers... recommend you revise basic calculus a little more as the last step is pretty basic maths... if it helps: a - x = b solve for x if you can do the above problem you should be able to clue out the final step to solve for h

OpenStudy (anonymous):

thanks bro .... have get rid of it